# Frequency-Dependent Squeezing for Gravitational-Wave Detection through Quantum Teleportation

### v2

@@ Abstract
Ground-based interferometric gravitational wave detectors are highly precise sensors for weak forces, limited in sensitivity across their detection band by quantum fluctuations of light. Current and future instruments address this limitation by injecting frequency-dependent squeezed vacuum into the detection port, utilizing narrow-band, low-loss optical cavities for optimal rotation of the squeezing ellipse at each signal frequency. This study introduces a novel scheme of such vacuum injection employing the principles of quantum teleportation. It allows achieving broadband suppression of quantum noise in detuned signal recycled-Fabry-Pérot–Michelson interferometers, which is the baseline design of the low-frequency detector within the Einstein Telescope xylophone detector, without requiring additional filter cavities or modifications to the core optics of the main interferometer.
@@ I Introduction
Figure 1: Schematics of frequency-dependent squeezing via quantum teleportation. We use two-photon formalism, where each beam is described by quadrature amplitude operators, where subscripts 1 and 2 denote amplitude and phase quadrature, respectively. All transformations of the beams are represented by transfer matrices. The Bell observables here are defined as $\hat{\alpha}_{1}=(\hat{V}_{1}-\hat{A}_{1})/\sqrt{2}$ and $\hat{\alpha}_{2}=(\hat{V}_{2}+\hat{A}_{2})/\sqrt{2}$. When Victor, Alice and Bob’s light undergo physical transformations, $\mathbb{U}_{v}$, $\mathbb{U}_{a}$ and $\mathbb{U}_{b}$, the teleported state is transformed accordingly to $\mathbb{U}_{b}\mathbb{U}_{a}\mathbb{U}_{v}\ket{\psi}$ accordingly.
@@
Gravitational wave detectors serving as highly precise displacement measurement instruments, are limited by quantum noise across most of their frequency band. At low frequencies, the optimal sensitivity of a conventional detector faces the constraints of the standard quantum limit (SQL) [16], as a natural consequence of Heisenberg’s uncertainty principle [17]. To overcome the SQL, scientists have proposed a variety of technologies based on the quantum non-demolition measurement principle [18, 19]. The mainstream approach for quantum noise reduction in the existing and future detectors is frequency-dependent squeezing (FDS) injection [20, 21]. The injected squeezed vacuum is produced by means of degenerate parametric down-conversion in nonlinear optical crystals. The squeezing angle at each frequency is optimized by passive optical cavities to cancel out the rotation of the squeezing ellipse due to ponderomotive effect within the interferometer.
@@
In tuned-broadband configurations, a single filter cavity is sufficient as indicated in prior studies [22, 23]. However, a detuned configuration, such as the low frequency interferometer of the Einstein Telescope’s xylophone detector (ETLF), potentially necessitates two filter cavities: one for compensating the optical detuning resonance in the interferometer, the other for compensating the so called optical spring resonance from the ponderomotive rigidity effect [16, 24]. These filter cavities, while presently at a hundred-meter scale in the second-generation detectors, are slated to reach kilometer scale in the third-generation observatories. The installation of kilometer-scale filter cavities entails substantial costs and technological difficulties. To tackle this challenge, several approaches to achieve the required sensitivity without the need for filter cavities have been proposed so far, such as the use of entangled light and negative-mass atomic spin ensembles [25, 26] and the small-scale optomechanical filters [27]. The EPR (for Einstein-Podolski-Rosen) or conditional squeezing has been proposed by Ma, et al. [28], and experimentally demonstrated in [29, 30, 31, 32]. In this scheme, by applying the concept of EPR steering [33, 34] to gravitational wave detectors, the main Fabry–Pérot-Michelson interferometer can be repurposed as a single filter cavity. However, such an EPR squeezing is insufficient for detuned configurations, which require two filter cavities.
@@
In this paper, we introduce the concept of quantum teleportation (QT) squeezing, which employs the principle of quantum teleportation to achieve frequency-dependent squeezing in a detuned dual-recycled Fabry–Pérot-Michelson interferometer without the use of filter cavities. The quantum teleportation of an optical state has been a well-established technique in the field of continuous-variable quantum information processing since it was first demonstrated by Furusawa et al. based on the Braustein–Kimble scheme  [35, 36]. The teleportation procedure in the Braustein–Kimble scheme involves Alice, the sending station; Bob, the receiving platform which shares EPR-entangled photons [37] and Victor, that brings an unknown quantum state and initiates the process (see Fig. 1). Alice conducts the Bell measurement involving Victor’s and her own photons, transmitting the outcome to Bob through the classical communication channel. Bob then displaces his photon based on Alice’s information, leading to the successful teleportation of Victor’s quantum state. By adding physical operations to Alice, Bob, and Victor’s paths, described as $\mathbb{U}_{a,b,v}$ in Fig 1, one can also manipulate the final teleported state. In our proposal of QT squeezing, we associate the main-interferometer response including the ponderomotive squeezing [20] with $\mathbb{U}_{b}$, and quadrature rotations by passive cavities with $\mathbb{U}_{v,a}$
@@
Fig. 2 shows the implementation of QT squeezing in a gravitational-wave detector. Three beams, Victor, Alice and Bob, are injected into the interferometer from the anti-symmetric port. The interferomter, acting as an empty cavity for Victor and Alice, provides an optimized frequency-dependent quadrature rotation by tuning their center frequencies and macroscopic lengths of the arm and signal extraction cavity, which functions as filter cavities. Bob’s frequency is matched to the main laser from the symmetric port, then he sees the interferometer as an active cavity including ponderomotive squeezing [20]. The output is spectrally separated into two detection ports by the output mode cleaners; Bob’s beam get collapsed via homodyne detection, while Victor and Alice are detected through Bell measurement. By applying Wiener filters to the two outputs of Bell measurement, quantum noise suppression is achieved as shown in Fig 4. In following sections, we will present the details of each step mentioned above.
@@
Figure 2: Configuration of QT squeezing. The OPA in the anti-symmetric port is pumped at two frequencies, $2\omega_{0}+\Delta_{a}$ and $2(\omega_{0}+\Delta_{v})$, generating entanglement at the sideband frequencies symmetrically (see the top-right box). The former pumping results in a two-mode EPR entanglement, Alice and Bob, while the latter forms a squeezed state, Victor. Three beams are injected through a Faraday isolator. The central part consists of a signal-recycled Fabry–Pérot Michelson Interferometer, which includes the beam splitter (BS), input-test mass (ITM), end-test mass (ETM) and signal-extraction mirror (SEM). The interferometer is pumped at the frequency $\omega_{0}$, matching Bob’s frequency. The output is spectrally separated by an output mode cleaner (OMC). Bob’s beam is collapsed at the homodyne detection, while Victor and Alice is detected through Bell measurement: two beams are combined with the local oscillator (LO) fields with the LO angle $\xi_{\mathrm{LO}}$, and subsequently detected by two photo detectors. The outputs are subtracted from each other and demodulated by two demodulation angles. The two sets of measurement data are combined using the optimal filter gain, $(g_{1}\ g_{2})$, and finally we achieve quantum-noise suppression.
@@ II Quantum state preparation
Figure 3: Schematics of three beams. Top and bottom panels show fields in the sideband and quadrature pictures, respectively. The OPA is pumped at two frequencies, $2\omega_{0}+\Delta_{a}$ and $2\omega_{0}+2\Delta_{v}$, generating entanglement at the sideband frequencies symmetrically. The former pumping results in a two-mode EPR entanglement, Alice and bob, while the latter forms a squeezed state, Victor.
The initial quantum states of three beams, Alice, Bob, Victor are prepared with a multi-mode squeezer [38]. The entangled beams, Alice and Bob, are centred at frequencies $\omega_{0}+\Delta_{a}$ and $\omega_{0}$, created by driving the OPA with a pumping beam at frequency $2\omega_{0}+\Delta_{a}$; the squeezed beam, Victor at frequency $\omega_{0}+\Delta_{v}$, is created by the second pumping beam at frequency $2(\omega_{0}+\Delta_{v})$ (see Fig. 3). Here $\Delta_{a}$ is supposed to be within the OPA’s bandwidth, and $\Delta_{v}-2\Delta_{a}$ equals to the multiples of the free spectral range of the OPA cavity.
@@ III.1 Bell measurement
The teleportation of Victor’s state to Bob requires a Bell measurement between Alice and Victor. In the Bell measurement, a local oscillator at frequency $\omega_{0}+(\Delta_{a}+\Delta_{v})/2$ is combined with Alice and Victor by a half beam splitter. Subsequently, the two output beams are detected by two photo detectors and their photon currents are demodulated at frequency $(\Delta_{v}-\Delta_{a})/2$. By setting the local oscillator angle to $-\pi/2$ and properly choosing the demodulation phase, Bell observables can be carried out (details are available in Appendix B):
@@ III.2 Post processing
$\begin{split}S^{\mathrm{tel}}_{\hat{B}_{2}\hat{B}_{2}}=|\Gamma|^{2}\frac{1+e^{-2r}\cosh 2r}{e^{-2r}+\cosh 2r}\xrightarrow{r\gg 1}|\Gamma|^{2}\frac{3}{e^{2r}}.\end{split}$ (9)
Taking into account that the gravitational-wave signal sidebands are not affected throughout the post-processing, Eq.(9) indicates that the strain sensitivity is improved by a factor of $3/e^{2r}$ across wide range of frequencies with sufficient squeezing.
@@ IV Sensitivity comparison
In the top panel of Fig 4, we show the quantum noise-limited sensitivity of ETLF [10] with the conventional frequency-dependent squeezing (hereafter called the baseline FDS) and QT squeezing. As a comparison we also plotted the enhancement factor compared to non-squeezed case of the baseline FDS (denoted as No SQZ baseline). The parameters considered, including imperfections such as losses and phase noises, are consistent with the current design of ETLF employing filter cavities, except for the squeezing level as shown in Table 1. Detuning frequencies $\Delta_{a}$ and $\Delta_{v}$ are also shown in the table, while the method of parameter searching and details of macroscopic length tuning are discussed in Appendix C.
@@
Detuning $\Delta_{a}$ $\sim 16.5$ MHz
Detuning $\Delta_{v}$ $\sim 1.59$ GHz
@@
However, the performance of QT squeezing can potentially surpass that of baseline FDS around the optical-spring resonance. The uneven sensitivity enhancement in the baseline scenarios arises from the dephasing of the squeezed vacuum. As studied in [43, 42], the dephasing of squeezed beam will lead coupling of the noise fluctuation from anti-squeezed quadrature into the squeezed quadrature. Around the resonance, where the effective mechanical susceptibility of test mass strengthens due to the optical spring effect, the interferometer creates more significant ponderomotive squeezing onto the quantum fields interacting with the mirrors. Therefore, the sensitivity is more susceptible to dephasing effect in particular at this frequency band, and the optimal input squeezing level is limited to $\sim\ 10$ dB (see the dashed and dotted curves in Fig 4). Such dephasing can stem from length fluctuation of optical paths and optical losses in detuned cavities. In the baseline FDS, the filter cavity length is constrained to 1 km due to infrastructure limitations, as also employed in [44]. In contrast, QT squeezing leverages the stability and length of the 10-km-long arm cavities as filter cavities. The length fluctuation of the arm cavities is well-suppressed by multi-stage suspension and control systems. In the particular case of ETLF with parameters in Table 1, QT squeezing exhibits better sensitivity than the baseline FDS at 8 Hz, as shown in Fig. 4.
As a filter cavity, the 10 km interferometer has lower effective loss compared with 1 km filter cavity, evaluated through the term loss per unit length [43]. We take the 1000 ppm loss from SEC into account, and it turns out the SEC loss is mitigated by the low transmissivity of the input-test masses (see Appendix D for more details). Those result in less effective loss and dephasing, thus allows higher input squeezing level. By accommodating -15 dB squeezing to compensate the 4.8 dB penalty, we can achieve $\sim$ 5 dB sensitivity improvement over the whole frequency band. In this particular case, higher squeezing level up to -17 dB allows to improve the sensitivity across the overall frequency range (see details in Fig. 9 in Appendix D).

### v3

@@ Abstract
Ground-based interferometric gravitational wave detectors are highly precise sensors for weak forces, limited in sensitivity across their detection band by quantum fluctuations of light. Current and future instruments address this limitation by injecting frequency-dependent squeezed vacuum into the detection port, utilizing narrow-band, low-loss optical filter cavities for optimal rotation of the squeezing ellipse at each signal frequency. This study introduces a novel scheme of such vacuum injection employing the principles of quantum teleportation, which works the same as arbitrary number of filter cavities without additional kilometer-scale infrastructure. We applied this scheme to detuned signal recycled-Fabry-Pérot–Michelson interferometers, which is the baseline design of the low-frequency detector within the Einstein Telescope xylophone detector. It is shown that our scheme achieves broadband suppression of quantum noise without requiring additional filter cavities or modifications to the core optics of the main interferometer.
@@ I Introduction
Figure 1: Schematics of a) EPR steering and b) quantum teleportation. We use two-photon formalism, where each beam is described by quadrature amplitude operators, where subscripts 1 and 2 denote amplitude and phase quadrature, respectively. All transformations of the beams are represented by transfer matrices, and $\pmb{\hat{a}}_{\mathrm{sqz}}$ is a squeezed vacuum. The Bell observables here are defined as $\hat{\alpha}_{1}=(\hat{V}_{1}-\hat{A}_{1})/\sqrt{2}$ and $\hat{\alpha}_{2}=(\hat{V}_{2}+\hat{A}_{2})/\sqrt{2}$. When Victor, Alice and Bob’s state undergo physical transformations, $\mathbb{U}_{v}$, $\mathbb{U}_{a}$, which are assumed to be phase rotations, and $\mathbb{U}_{b}$, which corresponds to the ponderomotive squeezing, the teleported state is transformed accordingly to $\mathbb{U}_{b}\mathbb{U}_{a}\mathbb{U}_{v}\pmb{\hat{v}}$ accordingly.
@@
Gravitational wave detectors serving as highly precise displacement measurement instruments, are limited by quantum noise across most of their frequency band. At low frequencies, the optimal sensitivity of a conventional detector faces the constraints of the standard quantum limit (SQL) [16], as a natural consequence of Heisenberg’s uncertainty principle [17]. To overcome the SQL, scientists have proposed a variety of technologies based on the quantum non-demolition measurement principle [18, 19]. The mainstream approach for quantum noise reduction in the existing and future detectors is frequency-dependent squeezing (FDS) injection [20, 21]. The injected squeezed vacuum is produced by means of degenerate parametric down-conversion in nonlinear optical crystals. The squeezing angle at each frequency is optimized by passive optical cavities to cancel out the rotation of the squeezing ellipse due to ponderomotive effect within the interferometer. The current gravitational-wave detectors employ a single filter cavity for their tuned-broadband configurations [22], as it is sufficient for broadband compensation of the ponderomotive squeezing angle in the narrow-band approximation [23, 24]. However, when this approximation breaks down, a larger number of filter cavities become necessary (see prior studies [25, 26] for example). In either case, as the scale of the detectors increases, these filter cavities, which are currently at the hundred-meter scale [27, 28], will need to reach the kilometer scale.
@@
The installation of kilometer-scale filter cavities entails substantial costs and technological difficulties. To tackle this challenge, several approaches have been proposed to achieve the required sensitivity without the need for filter cavities. These include the use of electromagnetically induced transparency [29], entangled light and negative-mass atomic spin ensembles [30, 31], and small-scale optomechanical filters [32]. The EPR (for Einstein-Podolski-Rosen) or conditional squeezing has been proposed by Ma, et al. [33], and experimentally demonstrated in [34, 35, 36, 37]. In this scheme, by applying the concept of EPR steering [38, 39] to gravitational wave detectors, the main Fabry–Pérot-Michelson interferometer can be repurposed as a single filter cavity (see Fig. 1a).
In this paper, we propose a generalized scheme of the EPR squeezing, employing the concept of quantum teleportation. The quantum teleportation of an optical state has been a well-established technique in the field of continuous-variable quantum information processing since it was first demonstrated by Furusawa et al. based on the Braustein–Kimble scheme  [40, 41]. Through our scheme, one can realize arbitrary number of operations of quadrature-angle rotation without the use of filter cavities.
@@
The teleportation procedure in the Braustein–Kimble scheme involves Alice, the sending station; Bob, the receiving platform which shares EPR-entangled photons [42] and Victor, that brings an unknown quantum state and initiates the process (see Fig. 1b). Alice conducts the Bell measurement involving Victor’s and her own photons, transmitting the outcome to Bob through the classical communication channel. Bob then displaces his photon based on Alice’s information, leading to the successful teleportation of Victor’s quantum state. By adding physical operations to Alice, Bob, and Victor’s paths, described as $\mathbb{U}_{a,b,v}$ in Fig 1b, one can also manipulate the final teleported state as $\mathbb{U}_{b}\mathbb{U}_{a}\mathbb{U}_{v}\pmb{\hat{v}}$. Furthermore, Bell measurement extends the limit of participating modes from 2 to an arbitrary number. By plugging the steered state or the teleported state to the initial state of Victor in a new teleportation protocol, $\pmb{\hat{v}}$, via Bell measurement, one can increase the number of equivalent filter cavities to an arbitrary number (see detailed discussion in Appendix B).
In the following sections, we illustrate how QT squeezing can be implemented in future detectors, which necessitates multiple filter cavities. Specifically, we focus on the low-frequency interferometer of the Einstein Telescope’s xylophone detector (ETLF). In the ETLF, the round-trip phase in the signal extraction cavity is slightly detuned to connect the suspended mirrors by the so-called optical spring [16, 43], which enables overcoming the Standard Quantum Limit (SQL). Additionally, for a narrow-band interferometer such as the ETLF, detuning offers a flexibility to boost the detector’s sensitivity [10].
@@
To realize broadband noise suppression for such a detuned configuration, two filter cavities are required: one is for compensating the optical detuning resonance in the interferometer, the other for compensating the so called optical spring resonance from the ponderomotive rigidity effect. We associate the main-interferometer response including the ponderomotive squeezing [20] with $\mathbb{U}_{b}$ in Fig 1b, and quadrature rotations by passive cavities with $\mathbb{U}_{v,a}$. Fig. 2 shows the implementation. Three beams, Victor, Alice and Bob, are injected into the interferometer from the anti-symmetric port. The interferometer, acting as an empty cavity for Victor and Alice, provides an optimized frequency-dependent quadrature rotation by tuning their center frequencies and macroscopic lengths of the arm and signal extraction cavity, which functions as filter cavities. Bob’s frequency is matched to the main laser from the symmetric port, then he sees the interferometer as an active cavity including ponderomotive squeezing [20]. The output is spectrally separated into two detection ports by the output mode cleaners; Bob’s beam get collapsed via homodyne detection, while Victor and Alice are detected through Bell measurement. By applying Wiener filters to the two outputs of Bell measurement, quantum noise suppression is achieved as shown in Fig 4. In following sections, we will present the details of each step mentioned above.
@@
Figure 2: Configuration of QT squeezing. The OPA in the anti-symmetric port is pumped at two frequencies, $2\omega_{0}+\Delta_{a}$ and $2(\omega_{0}+\Delta_{v})$, generating entanglement at the sideband frequencies symmetrically (see also Fig. 3b). The former pumping results in a two-mode EPR entanglement, Alice and Bob, while the latter forms a squeezed state, Victor. Three beams are injected through a Faraday isolator. The central part consists of a signal-recycled Fabry–Pérot Michelson Interferometer, which includes the beam splitter (BS), input-test mass (ITM), end-test mass (ETM) and signal-extraction mirror (SEM). The interferometer is pumped at the frequency $\omega_{0}$, matching Bob’s frequency. The output is spectrally separated by an output mode cleaner (OMC). Bob’s beam is collapsed at the homodyne detection, while Victor and Alice is detected through Bell measurement: two beams are combined with the local oscillator (LO) fields with the LO angle $\xi_{\mathrm{LO}}$, and subsequently detected by two photo detectors. The outputs are subtracted from each other and demodulated by two demodulation angles. The two sets of measurement data are combined using the optimal filter gain, $(g_{1}\ g_{2})$, and finally we achieve quantum-noise suppression.
@@ II Quantum state preparation
Figure 3: The OPA resonance for a) two pump fields and b) the signal and idlers. In this specific case two pump fields are separated by 5 FSR of the OPA cavity, and frequency difference between Victor and Bob is three time larger than that of Alice and Bob. c) Schematics of three beams. Top and bottom panels show fields in the sideband and quadrature pictures, respectively. The OPA is pumped at two frequencies, $2\omega_{0}+\Delta_{a}$ and $2\omega_{0}+2\Delta_{v}$, generating entanglement at the sideband frequencies symmetrically. The former pumping results in a two-mode EPR entanglement, Alice and Bob, while the latter forms a squeezed state, Victor.
The initial quantum states of three beams, Alice, Bob, Victor are prepared with a multi-mode squeezer [44]. The entangled beams, Alice and Bob, are centred at frequencies $\omega_{0}+\Delta_{a}$ and $\omega_{0}$, created by driving the OPA with a pumping beam at frequency $2\omega_{0}+\Delta_{a}$; the squeezed beam, Victor at frequency $\omega_{0}+\Delta_{v}$, is created by the second pumping beam at frequency $2(\omega_{0}+\Delta_{v})$ (see Fig. 3a). Here $\Delta_{a}$ is supposed to be a free spectral range (FSR) of the OPA cavity away from Bob’s frequency to keep resonant conditions of the entangled beam, and $\Delta_{a}-2\Delta_{v}$ equals to the multiples of the FSR of the OPA cavity (see Fig. 3).
@@ III.1 Bell measurement
The teleportation of Victor’s state to Bob requires a Bell measurement between Alice and Victor. In the Bell measurement, a local oscillator at frequency $\omega_{0}+(\Delta_{a}+\Delta_{v})/2$ is combined with Alice and Victor by a half beam splitter. Subsequently, the two output beams are detected by two photo detectors and their photon currents are demodulated at frequency $(\Delta_{v}-\Delta_{a})/2$. By setting the local oscillator angle to $-\pi/2$ and properly choosing the demodulation phase, Bell observables can be carried out (details are available in Appendix C):
@@ III.2 Post processing
$\begin{split}S^{\mathrm{tel}}_{\hat{B}_{2}\hat{B}_{2}}=|\Gamma|^{2}\frac{1+e^{-2r}\cosh 2r}{e^{-2r}+\cosh 2r}\xrightarrow{r\gg 1}|\Gamma|^{2}\frac{3}{e^{2r}}\end{split}$ (9)
(see detailed derivations in Appendix A). Taking into account that the gravitational-wave signal sidebands are not affected throughout the post-processing, Eq.(9) indicates that the strain sensitivity is improved by a factor of $3/e^{2r}$ in power across a wide range of frequencies with sufficient squeezing, which corresponds to 4.8 dB penalty discussed in Sec. V.
@@ IV Sensitivity comparison
In the top panel of Fig 4, we show the quantum noise-limited sensitivity of ETLF [10] with the conventional frequency-dependent squeezing (hereafter called the baseline FDS) and QT squeezing. As a comparison we also plotted the enhancement factor compared to non-squeezed case of the baseline FDS (denoted as No SQZ baseline). As shown, the detuned configuration creates a dip of optical spring at very-low frequency ($<10$ Hz), enabling to broaden the total detector’s bandwidth to lower frequency than a tuned configuration. The parameters considered, including imperfections such as losses and phase noises, are consistent with the current design of ETLF employing filter cavities, except for the squeezing level as shown in Table 1. Detuning frequencies $\Delta_{a}$ and $\Delta_{v}$ are also shown in the table, while the method of parameter searching and details of macroscopic length tuning are discussed in Appendix D.
@@
Detuning $\Delta_{a}$ $\sim 319$ MHz
Detuning $\Delta_{v}$ $\sim 962$ MHz
@@
However, the performance of QT squeezing can potentially surpass that of baseline FDS around the optical-spring resonance. The uneven sensitivity enhancement in the baseline scenarios arises from the dephasing of the squeezed vacuum. As studied in [49, 48], the dephasing of squeezed beam will lead coupling of the noise fluctuation from anti-squeezed quadrature into the squeezed quadrature. Around the resonance, where the effective mechanical susceptibility of test mass strengthens due to the optical spring effect, the interferometer creates more significant ponderomotive squeezing onto the quantum fields interacting with the mirrors and the dephasing from the filter cavity becomes large. Therefore, the sensitivity is more susceptible to dephasing effect in particular at this frequency band, and the optimal input squeezing level is limited to $\sim\ 10$ dB (see the dashed and dotted curves in Fig 4). Such dephasing can stem from length fluctuation of optical paths and optical losses in detuned cavities. In the baseline FDS, the filter cavity length is constrained to 1 km due to infrastructure limitations, as also employed in [50]. In contrast, QT squeezing leverages the stability and length of the 10-km-long arm cavities as filter cavities. The length fluctuation of the arm cavities is well-suppressed by multi-stage suspension and control systems. In the particular case of ETLF with parameters in Table 1, QT squeezing exhibits better sensitivity than the baseline FDS at 8 Hz, as shown in Fig. 4.
As a filter cavity, the 10 km interferometer has lower effective loss compared with 1 km filter cavity, evaluated through the term loss per unit length [49]. We take the 1000 ppm loss from SEC into account, and it turns out the SEC loss is mitigated by the low transmissivity of the input-test masses (see Appendix E for more details). Those result in less effective loss and dephasing, thus allows higher input squeezing level. By accommodating -15 dB squeezing to compensate the 4.8 dB penalty, we can achieve $\sim$ 5 dB sensitivity improvement over the whole frequency band. In this particular case, higher squeezing level up to -17 dB allows to improve the sensitivity across the overall frequency range (see details in Fig. 10 in Appendix E).